Kevin does not know the answer to two questions on a multiple choice exam.The first question has 5 choices,and the second question he does not know has 6 choices.What is the probability he will get both question wrong?
Answer by ilana(307)

The first question has 5 choices, so 4/5 are wrong. The second has 6 choices, so 5/6 are wrong. We can assume the accuracy of one has no affect on the accuracy of the other, so we have two independent probabilities and can multiply for the combined probability. (4/5)*(5/6)=(4*5)/(5*6)=20/30, or 2/3.